# The Astrometric Resoeccentric Degeneracy: Eccentric Single Planets Mimic 2:1 Resonant Planet Pairs in Astrometry

### v1

@@ 4.1 Exoplanet Astrometry in the Gaia Era
Gaia is projected to discover $\sim$1,900 long-period planets with the release of DR4 and $\sim$38,000 with DR5 (at <<20% precision; lammers2025exoplanet). Many more partial orbits and systems that cannot be as precisely constrained will also be observed (e.g., Wallace2025). While some giant planets on distant orbits are currently known, this dataset will represent our first population-level probe of the outer reaches of planetary systems and thus provides a unique opportunity to analyze the demographics and occurrence rates of long-period planets.
@@
Our work highlights the meticulous care that must be taken in interpreting these demographics. We predict that if single-planet solutions are assumed, an overabundance of eccentric planets may be reported. As we shall elaborate upon in §4.3, if these eccentric planets are incorrectly taken at face value, our understanding of their dynamical histories shall be led astray. If this degeneracy is not resolved by mutual inclination as discussed in §2.5, we encourage community efforts to pursue follow-up campaigns and continue to develop rigorous sampling techniques to address it, as outlined in §4.6.
@@
4.2 Long-Period Resonant Planets May Be Common
In multi-planet systems, interactions with the protoplanetary disk generally induce convergent inward migration (lin1986proto), making compact resonant chains the expected outcome of planetary formation (e.g. snellgrove2001disc; lee2002dynamics). Indeed, the outer Solar System is believed to have originated in such a resonant configuration (tsiganis2005; gomes2005; morbi2005). While dai2024resonance showed resonance is indeed common in young systems, this trend is not observed in more mature systems. The generally accepted picture is that resonant chains destabilized over the subsequent evolution of the system, a paradigm coined “breaking the chains” (izidoro2017chains).
@@ 4.2 Long-Period Resonant Planets May Be Common
4.3 Mutually Inclined Systems
The fact that mutual inclination eliminates this degeneracy (see §2.5) is fortuitous, given that the mutual inclination is also of paramount dynamical interest. Planets are expected to form in circular well-aligned orbits, as planet-disk interactions will damp out any eccentricity or inclination excursions (cresswell2007evolution). Hence, both eccentric planets and mutually inclined planets (e.g. xuan2020mutual; an2025significant; bgagliuffi2025jwst) may be signals of dynamically hot histories. However, the inferred evolutionary pathways differ for an isolated eccentric planet versus a system containing two misaligned planets. Below, we summarize the dynamical histories predicted by several mechanisms considered especially influential in sculpting exoplanetary architectures.
  • Planet-planet scattering (chatterjee_2008; juric2008dynamical; nagasawa2008scattering; ford2008; petrovich2014scattering; anderson2020insitu; marzari2025planet; lu2025hatp11) involves violent gravitational interactions induced by close encounters between pairs of planets, and can generate both significant eccentricity and inclination. raymond2008mmrs showed that while mean motion resonances are not the typical outcome of planet-planet scattering, this outcome does occur with appreciable frequency. They opine
@@ 4.3 Mutually Inclined Systems
    Planet-planet scattering events involving only two planets tend to eject one of them, and rarely generate extremely high mutual inclinations in the cases where both planets survive (ford2001dynamical; nagasawa2011orbital). Hence, a mutually inclined system likely originated with three or more planets while a single eccentric planet could be generated with only two planets, if scattering is responsible.
  • Secular interactions between planets can result in both excited eccentricities and inclinations (MurrayDermott1999), and these effects are greatly enhanced near mean motion resonances (e.g. malhotra1998resonances). If instead a single eccentric planet must be explained via secular interactions, the dynamical interpretation changes – the perturbing planet must be distant enough as to be undetectable. For instance, the system may be hierarchical and an unseen distant body may induce ZLK oscillations (vonzeipel1910zlk; lidov1962evolution; kozai1962secular; naoz2016eccentric).
  • Primordial misalignment. While eccentricity is very difficult to explain without invoking some sort of dynamical mechanism (e.g. duffell2015eccentric), the same cannot be said for misalignments. Primordial misalignments can arise naturally through processes that tilt the protoplanetary disk itself (lai2008wave; kuffmeier2021misaligned), a scenario for which there is growing observational support (BiddleBowler2025). Intriguingly, however, we note that such observational efforts are only capable of probing the outer reaches of protoplanetary disks – there is complementary evidence suggesting that inner disks are preferentially aligned (radzom2024misalignment; radzom2025misalignment). This supports the hypothesis that disks may often warp into two mutually misaligned sections (e.g. nixon2013tearing; nealon2018warp), resulting in mutually inclined planets.
@@
4.4 Peas-in-a-Pod-like Architecture
Our constraints on the mass ratios of distant giant planets are poor, both observationally and theoretically. A tempting assumption to make would be that of relatively equal-mass planets, a trend seen in the demographics of closer-in planets and coined “peas-in-a-pod” (millholland2017peas; weiss2018peas).
If indeed the outer reaches of planetary systems largely exhibit the same intra-system mass uniformity as their close-in counterparts, the degeneracy we have pointed out will not be a concern. Inspection of Equation 21 reveals that should the two planets in question be of equal mass, a single planet masquerading as the pair would require an unphysical eccentricity of $e_{\mathrm{eff}}$ = 1.26 to mimic the astrometric signal. More generally, in systems where $M_{p,2}\gtrsim 0.79M_{p,1}$ this degeneracy will not be a concern.
4.5 Hill-stability mass limits
The preferred parameter space of Gaia-detectable planets will be super-Jupiters on 25AU2-5\,AU orbits (lammers2025exoplanet). At first glance, this narrow semimajor-axis range may seem too restrictive to host multiple massive planets on long-term stable orbits. In this section, we demonstrate that such configurations are in fact readily achievable.
The Hill stability boundary of two planets on initially coplanar circular orbits is well understood. Gladman1993 derived an analytic criterion for the orbital spacing $\Delta\equiv a_{2}-a_{1}$ demanded for Hill stability for two planets of masses $M_{p,1}$ and $M_{p,2}$ orbiting a star of mass $M_{\star}$:
@@ 4.5 Hill-stability mass limits
We derive the maximum total planetary mass compatible with Hill stability for a pair of planets locked in a 2:1 period commensurability. From Kepler’s third law (kepler1609third), $\Delta\approx 0.587a_{1}$ for the 2:1 resonance. We may substitute this into Equation 24 and solve for the total planetary mass $M_{p,1}+M_{p,2}$:
@@
We note that Gladman1993 did not consider the stabilizing influence of mean motion resonances in their study. Numerous authors (e.g. wisdom1980resonance; marti2013gj876; obertas_2017; tamayo2017convergent; Tamayo20; lammers2024instability; lammers_winn_2024; lu_hip; tamayo2025unified) have demonstrated that exact mean motion resonances are associated with strong islands of stability. We present evidence in §4.2 that capture into these stable islands via convergent migration is an expected outcome of planet formation. This implies that larger planets may also be stable on long timescales in 2:1 configurations.
4.6 Other Ways to Break this Degeneracy
[…]
Combining radial velocity and astrometric observations is a powerful way to determine true planetary masses and orbital inclinations. Radial velocities measure only the line-of-sight motion and therefore yield the minimum mass, $M_{p}\sin i$, while astrometry measures the sky-plane component of the stellar reflex motion. When used together, these complementary data sets break the longstanding $\sin i$ degeneracy of radial velocity analyses (e.g., WrightHoward2009; Brandt2019; Winn2022; Yahalomi2023; Feng2024).
@@ 4.6 Other Ways to Break this Degeneracy
This RV-astrometry synergy has recently been demonstrated in practice via follow-up radial velocity monitoring to directly confirm and refine astrometric orbit solutions from Gaia DR3. In particular, RV follow-up of the Gaia-4b and Gaia-5b candidates revealed a massive planet and a brown dwarf orbiting low-mass stars, validating the astrometric orbits and providing precise dynamical masses (stefansson2025gaia). We note that even the Gaia-4b and Gaia-5b signals – with effective eccentricities $e_{\rm eff}\simeq 0.34$ and $e_{\rm eff}\simeq 0.64$, respectively, from the single-planet fits in stefansson2025gaia – could in principle be mimicked by circular 2:1 planet pairs: for Gaia-4, a pair with periods $\{571,286\}\,\mathrm{days}$ and masses $\{11.8,3.2\}\,M_{\rm Jup}$; and for Gaia-5, a pair with periods $\{359,179\}\,\mathrm{days}$ and masses $\{20.9,10.6\}\,M_{\rm Jup}$. While a detailed investigation of the resoeccentric degeneracy for these two systems is beyond the scope of this work, such an analysis would help clarify whether, and under what conditions, RV follow-up of astrometrically identified systems can break this degeneracy. Future work should additionally investigate whether RV breaking of the resoeccentric degeneracy is feasible for the types of planets we expect to find in Gaia DR4 and DR5 astrometry based on current RV capabilities and study the optimal orbital strategies for such observations.
@@
This work predicts an additional degeneracy that may appear in the Gaia epoch astrometric data. It is already well-documented that astrometric solutions become highly degenerate when the orbital period extends beyond the observational baseline (e.g., Casertano2008). This highlights the need for the development of proper astrometric inference techniques that are designed to explore highly degenerate and multi-modal parameter spaces. Modern orbit-fitting frameworks such as The Joker (thejoker), orbitize! (orbitize1; orbitize2), OFTI (ofti), Octofitter (octofitter), and orvara (orvara) have demonstrated the importance of using samplers that remain robust in the presence of strong covariances, long-lived degeneracies, and disconnected orbital solutions. Techniques such as rejection sampling, parallel tempering, and nested sampling have been developed precisely for this purpose, providing reliable posterior exploration even when the likelihood surface contains multiple isolated modes or narrow curved manifolds. These approaches are essential for accurately capturing the full family of orbital solutions implied by astrometric and radial-velocity data, and we strongly encourage their further development and study in preparation for Gaia DR4 and DR5. It will be important to understand in which regions of parameter space these different sampling methods optimally sample the complete solution space.
@@
We also encourage future work to investigate the signal-to-noise ratio (SNR) required to differentiate between the 2:1 and eccentric configurations. This will likely be a function of the masses, periods, and other orbital parameters (such as eccentricity and mutual inclination of the 2:1 planets), but will be a powerful result in ensuring over what regions of parameter space are statistically differentiated solutions possible. See kippingRVsnr2024 for the analogous derivation of this for RV data, which was built on the inclusion of higher order eccentricity terms as derived in Lucy2005.
@@
Finally, it may be possible to detect this resoeccentric degeneracy on a population level by searching for non-isotropic solutions of $\omega$ (the argument of the periastron) of single eccentric planets fit to the Gaia data. As the $\omega$ term would need to be “fine-tuned” in order to mimic 2:1 astrometric signals, anisotropies in the resulting $\omega$ distribution are likely – one could potentially flag that this effect is present at the population level. Injection-recovery of 2:1 systems fit with only single eccentric solutions could potentially confirm this degeneracy-flagging potential before Gaia DR4.
5 Conclusion

### v2

@@ IV.1 Exoplanet Astrometry in the Gaia Era
Gaia is projected to discover $\sim$1,900 long-period planets with the release of DR4 and $\sim$38,000 with DR5 (at <<20% mass and period precision; Lammers and Winn, 2025), predominantly at orbital separations around $2-5\,\mathrm{AU}$. Many more partial orbits and systems that cannot be as precisely constrained will also be observed (e.g., Wallace et al., 2025) at even further orbital separations. While some giant planets on multi-AU orbits are currently known, this dataset will represent the first population-level probe of planetary systems at several astronomical units, providing a unique opportunity for analysis of demographics and occurrence rates of these long-period planets.
@@
Our work highlights the meticulous care that must be taken in interpreting these demographics. We predict that if single planet solutions are assumed, an overabundance of eccentric planets may be reported. As we shall elaborate upon in §IV.3, if these eccentric planets are incorrectly taken at face value, our understanding of their dynamical histories shall be led astray. If this degeneracy is not resolved by mutual inclination as discussed in §II.5, we encourage community efforts to pursue follow-up campaigns and continue to develop rigorous sampling techniques to address it, as outlined in §IV.6.
@@
IV.2 Long-Period Resonant Planets May Be Common
In multi-planet systems, interactions with the protoplanetary disk generally induce convergent inward migration (Lin and Papaloizou, 1986), making compact resonant chains the expected outcome of planetary formation (e.g. Snellgrove et al., 2001; Lee and Peale, 2002). Indeed, the outer Solar System is believed to have originated in such a resonant configuration (Tsiganis et al., 2005; Gomes et al., 2005; Morbidelli et al., 2005). While Dai et al. (2024) showed resonance is indeed common in young systems, this trend is not observed in more mature systems. The generally accepted picture is that resonant chains destabilized over the subsequent evolution of the system, a paradigm coined “breaking the chains” (Izidoro et al., 2017).
@@ IV.2 Long-Period Resonant Planets May Be Common
We note that migration of giant planets that have been captured into mean-motion resonance leads to the excitation of eccentricity, potentially to high values (e.g. Lee and Peale, 2002). The non-zero but moderate eccentricities typically found in known systems of resonant giant planets (e.g. Lee and Peale, 2002; Correia et al., 2009; Rivera et al., 2010) are consistent with this eccentricity growth mechanism being balanced against damping from planet-disk interactions (Papaloizou and Larwood, 2000; Cresswell et al., 2007). For the high mass systems that will dominate the Gaia population, simulations and analytic arguments show that the strength of damping due to the disk interaction may be substantially weaker, or even switch sign to become eccentricity growth (Papaloizou et al., 2001; Goldreich and Sari, 2003; Ogilvie and Lubow, 2003). This could lead to substantially larger eccentricities for higher mass resonant pairs. A detailed analysis of the eccentricity limits in which this degeneracy holds is left to future work (see §IV.6 for more discussion). We note that these forthcoming Gaia observations are expected to significantly improve population-level constraints on the eccentricity distribution of resonant giant planets, motivating consideration of a broad range of dynamical outcomes.
If many systems are discovered in Gaia astrometry with very massive planets in 2:1 configurations, it would suggest that these planets likely formed in the disk and were then captured into resonance – thus implying a very massive disk. This would be an interesting result in its own right and could potentially offer indirect constraints on planetary formation models.
IV.3 Mutually Inclined Systems
The fact that mutual inclination can eliminate this degeneracy (see §II.5 and Figure 2) is perhaps fortuitous, given that the mutual inclination is also of paramount dynamical interest. Planets are expected to form in circular well-aligned orbits, as planet-disk interactions will damp out any eccentricity or inclination excursions (Cresswell et al., 2007). Hence, both eccentric planets and mutually inclined planets (e.g. Xuan and Wyatt, 2020; An et al., 2025; Bardalez Gagliuffi et al., 2025) may be signals of dynamically hot histories. However, the inferred evolutionary pathways differ for an isolated eccentric planet versus a system containing two misaligned planets. Below, we summarize the dynamical histories predicted by several mechanisms considered especially influential in sculpting exoplanetary architectures.
  • Planet-planet scattering (Chatterjee et al., 2008; Jurić and Tremaine, 2008; Nagasawa et al., 2008; Ford and Rasio, 2008; Petrovich et al., 2014; Anderson et al., 2020; Marzari, 2025; Lu et al., 2025a) involves violent gravitational interactions induced by close encounters between pairs of planets, and can generate both significant eccentricity and inclination. Raymond et al. (2008) showed that while mean motion resonances are not the typical outcome of planet-planet scattering, this outcome does occur with appreciable frequency. They state
@@ IV.3 Mutually Inclined Systems
    Planet-planet scattering events involving only two planets tend to eject one of them, and rarely generate extremely high mutual inclinations in the cases where both planets survive (Ford et al., 2001; Nagasawa and Ida, 2011). Hence, a mutually inclined system likely originated with three or more planets while a single eccentric planet could be generated with only two planets, if scattering is responsible.
  • Secular interactions between planets can result in both excited eccentricities and inclinations (Murray and Dermott, 1999), and these effects are greatly enhanced near mean motion resonances (e.g. Malhotra, 1998). If instead a single eccentric planet must be explained via secular interactions, the dynamical interpretation changes – the perturbing planet must be distant enough as to be undetectable. For instance, the system may be hierarchical and an unseen distant body may induce ZLK oscillations (von Zeipel, 1910; Lidov, 1962; Kozai, 1962; Naoz, 2016).
  • Primordial misalignment. While eccentricity is very difficult to explain without invoking some sort of dynamical mechanism (e.g. Duffell and Chiang, 2015), the same cannot be said for misalignments. Primordial misalignments can arise naturally through processes that tilt the protoplanetary disk itself (Lai and Zhang, 2008; Kuffmeier et al., 2021), a scenario for which there is growing observational support (Biddle et al., 2025). Intriguingly, however, we note that such observational efforts are only capable of probing the outer reaches of protoplanetary disks – there is complementary evidence suggesting that inner disks are preferentially aligned (Radzom et al., 2024, 2025). This supports the hypothesis that disks may often warp into two mutually misaligned sections (e.g. Nixon et al., 2013; Nealon et al., 2018), resulting in mutually inclined planets.
@@
IV.4 Peas-in-a-Pod-like Architecture
Our constraints on the mass ratios of distant giant planets are poor, both observationally and theoretically. A tempting assumption to make would be that of relatively equal-mass planets, a trend seen in the demographics of closer-in planets and coined “peas-in-a-pod” (Millholland et al., 2017; Weiss et al., 2018).
If indeed the outer reaches of planetary systems largely exhibit the same intra-system mass uniformity as their close-in counterparts, the degeneracy we have pointed out will not be a concern. Inspection of Equation 21 reveals that should the two planets in question be of equal mass, a single planet masquerading as the pair would require an unphysical eccentricity of $e_{\mathrm{eff}}$ = 1.26 to mimic the astrometric signal. More generally, in systems where $M_{p,2}\gtrsim 0.79M_{p,1}$ this degeneracy will not be a concern – and in fact higher order terms in eccentricity will likely break the resoeccentric degeneracy at even smaller mass ratios.
IV.5 Hill-stability mass limits
The preferred parameter space of Gaia-detectable planets will be super-Jupiters on 25AU2-5\,AU orbits (Lammers and Winn, 2025). At first glance, this narrow semimajor-axis range may seem too restrictive to host multiple massive planets on long-term stable orbits. In this section, we demonstrate that such configurations are in fact readily achievable.
The Hill stability boundary of two planets on initially coplanar circular orbits is well understood. Gladman (1993) derived an analytic criterion for the orbital spacing $\Delta\equiv a_{2}-a_{1}$ demanded for Hill stability for two planets of masses $M_{p,1}$ and $M_{p,2}$ orbiting a star of mass $M_{\star}$:
@@ IV.5 Hill-stability mass limits
We derive the maximum total planetary mass compatible with Hill stability for a pair of planets locked in a 2:1 period commensurability. From Kepler’s third law (Kepler, 1609), $\Delta\approx 0.587a_{1}$ for the 2:1 resonance. We may substitute this into Equation 24 and solve for the total planetary mass $M_{p,1}+M_{p,2}$:
@@
We note that Gladman (1993) did not consider the stabilizing influence of mean motion resonances in their study. Numerous authors (e.g. Wisdom, 1980; Martí et al., 2013; Obertas et al., 2017; Tamayo et al., 2017, 2020; Lammers et al., 2024; Lammers and Winn, 2024; Lu et al., 2025b; Tamayo and Hadden, 2025) have demonstrated that exact mean motion resonances are associated with strong islands of stability. We present evidence in §IV.2 that capture into these stable islands via convergent migration is an expected outcome of planet formation. Additionally, it is worth noting that for giant planets, the formation of resonances is likely to occur while the protoplanetary disk is present. In this case, tidal damping adds an extra stabilizing factor not considered in the analysis by Gladman (1993). Altogether, this implies that even high-mass planets may also be stable on long timescales in 2:1 configurations.
IV.6 Other Ways to Break this Degeneracy
The mathematical derivations in §II are first order approximations for the single eccentric planetary case and assume circular and coplanar orbits for the 2:1 resonant case. There is thus some eccentricity and mutual inclination at which this approximation will no longer hold. This will depend on a number of factors, including: orbital parameters (most dominantly masses and periods) of the planet(s) in the system; observational limitations, which will also depend on the distance to, location of, and brightness of the host star; and modeling choices. Additionally, in these derivations, we have assumed Keplerian orbits. As such, we leave an in depth analysis of the eccentricities and mutual inclinations at which the degeneracy is likely to be broken for systems in Gaia DR4 and DR5 to future work. In what follows, we discuss some possible strategies for mitigating and properly accounting for the resoeccentric degeneracy in Gaia astrometry.
A transit observation could break this degeneracy, as an additional constraint on the eccentricity can be derived from the transit alone via the photo-eccentric effect (Kipping et al., 2012; Dawson and Johnson, 2012). However, as transits are more sensitive to close-orbiting planets, we anticipate a minimal overlap between the population of Gaia astrometric planets and systems with observed transits. We can estimate the number of Gaia astrometric planets to transit by adopting the $\sim$1,900 planets in DR4 and $\sim$38,000 planets in DR5 with 20% mass and period precision from Lammers and Winn (2024). We use the planet transit probability formula assuming a circular orbit: $\mathcal{P}_{\text{tr}}=\frac{R_{\star}+R_{p}}{a}$ where $R_{\star}$ is the stellar radius, $R_{p}$ is the planetary radius, and $a$ is the semi-major axis. For a Jupiter radius planet orbiting a solar radius star at 25AU2-5\,AU, this equates to a transit probability of $\sim 0.21\%-0.084\%$. For simplicity, we adopt the mean transit percentage ($\sim 0.15\%)$ and predict that $\sim$3 Gaia DR4 planets will transit and $\sim$57 Gaia DR5 planets will transit. Any such systems will be of particular interest to better understand this degeneracy in Gaia astrometry, and may require targeted follow-up to obtain these transit observations.
Significant effort has been dedicated to disentangling this resoeccentric degeneracy in radial velocity observations. As previously noted, Anglada-Escudé et al. (2010) originally identified that $\sim$35% of single eccentric RV planets are statistically compatible with circular two-planet 2:1 configurations. Wittenmyer et al. (2012) presented an additional planet in this 2:1 configuration in two system previously identified as single eccentric planet systems. Wittenmyer et al. (2013) investigated 82 known (at the time) moderately eccentric single planet systems ($e>0.3$) to test the hypothesis that some may actually be low-eccentricity two-planet systems – identifying 9 particular systems of interest. A Bayesian model comparison of 60 moderately eccentric single planet systems with RV data identified 15 systems that favor 2:1 resonant near-circular orbits at 95% confidence (Boisvert et al., 2018). Wittenmyer et al. (2019) showed that RV-observed systems with two near-resonant planets on near-circular orbits most often masquerade as a single planet system with orbital eccentricity of approximately $e=0.31\pm 0.12$ in sparsely sampled data and that such systems were highly unlikely to imitate single planet systems with orbital eccentricities greater than $e\sim 0.5$. Presumably there will be a similar eccentricity at which this degeneracy will be most impactful and a maximum eccentricity at which this degeneracy will break in Gaia astrometry. We leave an investigation into these eccentricity limits to future work.
@@ IV.6 Other Ways to Break this Degeneracy
Combining radial velocity and astrometric observations is a powerful way to determine true planetary masses and orbital inclinations. Radial velocities measure only the line-of-sight motion and therefore yield the minimum mass, $M_{p}\sin i$, while astrometry measures the sky-plane component of the stellar reflex motion. When used together, these complementary data sets break the longstanding $\sin i$ degeneracy of radial velocity analyses (e.g., Wright and Howard, 2009; Brandt et al., 2019; Winn, 2022; Yahalomi et al., 2023; Feng, 2024). This RV-astrometry synergy has recently been demonstrated in practice via follow-up radial velocity monitoring to directly confirm and refine astrometric orbit solutions from Gaia DR3. In particular, RV follow-up of the Gaia-4b and Gaia-5b candidates revealed a massive planet and a brown dwarf orbiting low-mass stars, validating the astrometric orbits and providing precise dynamical masses (Stefánsson et al., 2025). We note that even the Gaia-4b and Gaia-5b signals – with effective eccentricities $e_{\rm eff}\simeq 0.34$ and $e_{\rm eff}\simeq 0.64$, respectively, from the single planet fits in Stefánsson et al. (2025) – could in principle be mimicked by circular 2:1 planet pairs: for Gaia-4, a pair with periods $\{571,286\}\,\mathrm{days}$ and masses $\{11.8,3.2\}\,M_{\rm Jup}$; and for Gaia-5, a pair with periods $\{359,179\}\,\mathrm{days}$ and masses $\{20.9,10.6\}\,M_{\rm Jup}$. While a detailed investigation of the resoeccentric degeneracy for these two systems is beyond the scope of this work, such an analysis would help clarify whether, and under what conditions, RV follow-up of astrometrically identified systems can break this degeneracy. Future work should additionally investigate whether RV breaking of the resoeccentric degeneracy is feasible for the types of planets we expect to find in Gaia DR4 and DR5 astrometry based on current RV capabilities and study the optimal orbital strategies for such observations.
@@
This work predicts an additional degeneracy that may appear in the Gaia epoch astrometric data. It is already well-documented that astrometric solutions become highly degenerate when the orbital period extends beyond the observational baseline (e.g., Casertano et al., 2008). This highlights the need for the development of proper astrometric inference techniques that are designed to explore highly degenerate and multi-modal parameter spaces. Modern orbit-fitting frameworks such as The Joker (Price-Whelan et al., 2017), orbitize! (Blunt et al., 2020, 2024), OFTI (Blunt et al., 2017), Octofitter (Thompson et al., 2023), and orvara (Brandt et al., 2021) have demonstrated the importance of using samplers that remain robust in the presence of strong covariances, long-lived degeneracies, and disconnected orbital solutions. Techniques such as rejection sampling, parallel tempering, and nested sampling have been developed precisely for this purpose, providing reliable posterior exploration even when the likelihood surface contains multiple isolated modes or narrow curved manifolds. These approaches are essential for accurately capturing the full family of orbital solutions implied by astrometric and radial-velocity data, and we strongly encourage their further development and study in preparation for Gaia DR4 and DR5. It will be important to understand in which regions of parameter space these different sampling methods optimally sample the complete solution space.
@@
We also encourage future work to investigate the signal-to-noise ratio (SNR) required to differentiate between the 2:1 and eccentric configurations. This will likely be a function of the masses, periods, and other orbital parameters (such as eccentricity and mutual inclination of the 2:1 planets), but will be a powerful result in ensuring over what regions of parameter space are statistically differentiated solutions possible. See Kipping and Wang (2024) for the analogous derivation of this for RV data, which was built on the inclusion of higher order eccentricity terms as derived in Lucy (2005).
@@
Finally, it may be possible to detect this resoeccentric degeneracy at the population level by searching for non-isotropic solutions of $\omega$ (the argument of the periastron) of single eccentric planets fit to the Gaia data. As the $\omega$ term would need to be “fine-tuned” in order to mimic the astometric signal of a 2:1 planetary configuration, the presence of unresolved resonant pairs would be expected to induce anisotropis in the inferred $\omega$ distribution. Such anisotropies could therefore serve as a population-level indicator of this degeneracy. Injection-recovery tests of 2:1 systems fit with single eccentric solutions could be used to validate this degeneracy-flagging approach prior to Gaia DR4.
V Conclusion